All That Blooms provides environmentally friendly lawn services for homeowners. Its operating costs are as follows.

Depreciation $1,224 per month
Advertising $153 per month
Insurance $1,683 per month
Weed and feed materials $10 per lawn
Direct labor $8 per lawn
Fuel $2 per lawn


All That Blooms charges $50 per treatment for the average single-family lawn.

Determine the company
Answer #1

Total fixed costs = depreciation + advertising + insurance = 1224 + 153 + 1683 = $3060/month

Total variable cost per unit = weed + direct labor + fuel = 10 + 8 + 2 = $20/lawn

Contribution margin ratio = (sales per unit - variable cost per unit) / sales per unit = (50 - 20) / 50 = 60%

Breakeven sales = fixed costs / contribution margin ratio = 3060 / 60% = $ 5100

Breakeven sales units = breakeven sales / sales per unit = 5100 / 50 = 102 lawns

Answer: Breakeven point = 102 lawns and breakeven sales point = $5100

Hope this helped ! Let me know in case of any queries.